OpenStudy (freethinker):

@ohyeahh I am posting your problem for you. The numbers in each row and column of the square are written according to the same rule. By defining the rule, find the missing number. 5 | 11 | 36 ------------- 11 | 4 | 125 ------------- 15 | 15 | ?

OpenStudy (freethinker):

36 - 11 = 25; 25 is 5^2 or you can say that 5 is sqrt(25) ------------------------- 125 - 4 = 121; 121 is 11^2 or you can say 11 is sqrt(121) ------------------------- x - 15 = 15^2 x - 15 = 225 +15 +15 x = 240 or you can do it with the sqrt sqrt(x-15) = 15 // square both sides x - 15 = 225 and you do the same algebraic manipulation + 15 +15 x = 240 is 240 a given? http://prntscr.com/14jd7p your choices are: a) 240 b) 250 c) 290 d) 300

OpenStudy (anonymous):

*facepalm* i should have got it.. thank you. :) but umm.. thats not the right answer. :o

OpenStudy (freethinker):

it is not? lol

OpenStudy (anonymous):

I mean it looks right.. but the book says its not.

OpenStudy (freethinker):

so what was the book's explanation or rule?

OpenStudy (anonymous):

It doesnt give any explanation. It just gives the correct answer. Its 250.

OpenStudy (freethinker):

interesting!

OpenStudy (freethinker):

then let us look for another rule that will satisfy that answer and will will be applicable to the rest

OpenStudy (aravindg):

240 is an answer .Is this question multiple answer type?

OpenStudy (anonymous):

Nope. single answer type.

OpenStudy (aravindg):

Then it might be a printing mistake

OpenStudy (anonymous):

hmm i guess. 240 looks right.

OpenStudy (anonymous):

Ok so, 240 is right most probably. Thank you. :)

OpenStudy (freethinker):

let us not give up on the idea that it might be a different answer.

OpenStudy (anonymous):

hmm yeah.